Ten grams of nitrogen is mixed with 5 grams of O2 and held at 25 degree Celcius at 0.750 bar.

a) What are the mole fractions of N2 and O2?
b) What are the partial pressures of N2 and O2?
c) What is the molar volume assuming the molecules do not interact?
d) What is the actual volume assuming the molecules do not interact?
1
Expert's answer
2020-12-06T13:51:28-0500

a)      Moles of N2 =10/28.02= 0.3569moles

Moles of O2= 5/16= 0.3125 Moles

Total number of moles of two gases= (0.3569+0.3125) moles

                                                                     =0.6694moles

Mole fraction of N2= 0.3569/0.6694= 0.5332

Moles of O2= 0.3125/0.6694= 0.4668


b)     Partial pressure of N2=0.5332×0.740192= 0.3947atm

[According to online conversion, 0.74bars=0.740192 atm]

Partial pressure of O2= 0.4668×0.740192=0.3454 atm

c)      1 molar volume of the two gases at the prevailing conditions is 24.00L

PV=nRT

V=nRT/P

    n=0.6694moles

P= (0.3947+0.3454) =0.7401atm

V=0.6694×0.0821×298/0.7401

 =22.13L